On the Apollo 14 mission to the moon, astronaut Alan Shepard hit a golf ball with a 6 iron. The acceleration due to gravity on the moon is 1/6 of its value on earth.Suppose he hits the ball with a speed of 17 m/s at an angle 60 above the horizontal.

How much farther did the ball travel on the moon than it would have on earth?

For how much more time was the ball in flight?
Answer #2

v is the velocity of the ball
v1 is the vertical velocity
v2 is the horizontal velocity

v1=v cos 30
v1= 14.72

v2=v sin30
v2= 8.5

time of flight on moon = v1/gm(where gm = 10/6)
= 8.832 sec

horizontal distance travelled =8.832 * v2= 75.07 m

time on earth would have been=v1/g(where g = 10)

=1.472

horizontal distance travelled on earth =1.472 * v2= 12.512 m

extra distance travelled = 62.56 m

extra time of flight = 7.36 sec

answered by: miguel
Answer #1
convert the speed into x y direction.
x: 17 cos 60 = 8.5 m/s
y: 17 sin 60 = 14.72 m/s

gravity: 1/6 * 9.81 = 1.635 m/s^2

find the time first:
time taken on moon = (0 - 14.72)/-1.635 = 9 seconds
time taken on earth = (0 - 14.72)/-9.81 = 1.5 seconds
difference = 9 - 1.5 = 7.5 seconds

distance travel on moon = 0.5 (0 + 8.5) * 9 = 38.25 m
distance travel on moon = 0.5 (0 + 8.5) * 1.5 = 6.375 m
difference = 38.25 - 6.375 = 31.875 m
answered by: The Curse Word Destroyer, age 11